### 05 Dec 2023 v2

# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

Hulk: A Universal Knowledge Translator
for Human-Centric Tasks
Yizhou Wang*, Yixuan Wu*, Shixiang Tang, Weizhen He, Xun Guo, Feng Zhu, Lei Bai, Rui Zhao,
Jian Wu,  Tong He, Wanli Ouyang
Yizhou Wang and Yixuan Wu are contributed equally. This work was done when Yixuan Wu, Xun Guo were interns at Shanghai AI Laboratory, Shanghai, China. Yixuan Wu is with the School of Medicine, Zhejiang University, Hangzhou, 310027, China. Weizhen He is with the College of Electrical Engineering, Zhejiang University, Hangzhou, 310027, China. Xun Guo is with the Department of Automation, University of Science and Technology of China, Hefei, 230052, China. Shixiang Tang is with the School of Electrical and Information Engineering, University of Sydney, NSW, Australia. Feng Zhu and Rui Zhao are with SenseTime Group Limited, China. Jian Wu is with Second Affiliated Hospital School of Medicine, and School of Public Health, Zhejiang University, Hangzhou, 310058, China. Yizhou Wang, Lei Bai, Tong He, and Wanli Ouyang are with Shanghai AI Laboratory, Shanghai, 200232, China.Manuscript received April 19, 2021; revised August 16, 2021.
I Full Ablation results
As shown in Table I, we present the full experimental results of the ablation studies. Specifically, we explore the affects of (1) different weight sharing strategies, (2) attention mask designs in decoder, (3) task collaboration and interference, and (4) different initialization weights. One can see that the same conclusion can be observed that as more model components are shared, Hulk can achieve higher performance. Besides, omitting any task generally leads to reduced performance. Furthermore, employing different attention interaction methods for different tasks can optimize the model’s performance, compared with using pure full attention or diagnal attention. Last, using the pre-trained weights from other human-centric models (i.e., HAP [yuan2023hap] and PATH [tang2023humanbench]) can further boost our hulk’s performance. This result not only validates the robustness of Hulk but also highlight its potential for further enhancement.
Moreover, in Table II, to assess the impact of the diverse and sufficient human-centric data, we train Hulk using a 60000-iteration schedule on the ablation datasets comprising 1.7M training samples, and full datasets comprising 30M training samples, respectively. Results indicate that incorporating more human-centric data and increasing the training iterations brings considerable performance improvement across all tasks.
TABLE I: Complete experimental results of different ablation studies.
Methods Parsing 2D Pose Detection Attribute Caption Skeleton 3D Pose & Mesh Avg.
H3.6 LIP CIHP COCO AIC Crowd PA rapv2 CUHK ntu60 Human3.6M 3DPW
mIoU mIoU mIoU AP AP AP mA mA B@4 Acc 100-MPJPE 100-PA-MPJPE 100-MPVPE 100-MPJPE 100-PA-MPJPE
(a) baseline (Hulk) 62.4 56.8 59.7 70.8 25.5 81.8 79.4 77.9 28.0 93.2 45.9 60.5 -0.3 13.2 49.0 53.6
Weight sharing
(b) only enc./dec. shared 63.4 56.8 59.6 71.1 25.8 80.6 80.8 78.8 26.7 93.8 45.2 60.3 -3.0 10.7 47.0 53.2
(c) only enc. shared 51.6 56.3 60.5 71.3 25.7 75.9 82.2 81.3 28.7 1.7 35.9 56.5 -11.2 2.9 41.3 44.0
Attention Mask Designs in the Decoder
full-attn 62.6 57.2 59.9 70.9 25.4 79.9 79.2 77.6 28.6 91.8 44.3 60.1 -1.0 13.4 48.1 53.2
diag-attn 62.7 56.6 59.9 70.6 25.6 80.3 79.3 78.0 28.4 93.1 40.3 57.6 -4.1 10.0 46.3 52.3
Task Collaboration and Interference
w/o attribute 62.4 56.9 59.9 70.7 25.6 79.9 28.2 93.1 44.4 60.0 -2.1 12.2 47.9 42.6
w/o caption 62.7 57.3 59.9 70.7 25.6 80.2 79.3 77.8 92.9 46.5 61.3 -2.2 11.2 47.5 51.4
w/o skeleton 62.3 57.4 60.1 70.7 25.7 80.9 79.1 77.4 28.2 45.7 59.6 -4.8 10.0 46.7 46.6
w/o mesh&3d pose 62.6 57.4 60.3 70.8 25.5 81.0 79.8 78.0 27.9 93.1 75.8
w/o detection 63.2 57.6 60.9 71.4 26.4 79.9 77.6 27.8 93.1 46.0 59.8 -4.8 10.5 47.6 47.8
w/o 2d pose 56.6 52.3 53.0 65.0 75.3 75.1 26.6 93.5 24.5 48.4 -18.9 -3.5 37.2 39.0
w/o parsing 70.2 25.2 78.5 79.4 77.3 25.8 93.2 42.5 57.5 -2.9 10.8 46.7 40.3
Different initialization weights
MAE (baseline) 62.4 56.8 59.7 70.8 25.5 81.8 79.4 77.9 28.0 93.2 45.9 60.5 -0.3 13.2 49.0 53.6
HAP 65.8 63.3 68.8 76.5 33.4 85.4 85.5 81.8 29.6 91.5 55.2 66.6 9.6 22.9 55.6 59.4
PATH 66.3 63.6 69.2 76.8 34.5 88.6 83.3 80.2 29.9 92.8 54.6 66.7 8.3 21.3 54.5 59.4
TABLE II: Complete experimental results of our Hulk performance on diverse human-centric tasks when scaling up data size. Experiments are both trained with a FULL 60,000-iteration schedule, showing that learn from additional data and achieve better performance on average. .
Methods Parsing 2D Pose Detection Attribute Caption Skeleton 3D Pose & Mesh Avg
H3.6 LIP CIHP COCO AIC Crowd PA rapv2 CUHK ntu60 Human3.6M 3DPW
mIoU mIoU mIoU AP AP AP mA mA B@4 Acc 100-MPJPE 100-PA-MPJPE 100-MPVPE 100-MPJPE 100-PA-MPJPE
30M 68.1 64.0 70.6 77.0 34.5 90.7 82.9 80.9 31.1 93.8 56.4 68.1 20.2 33.0 60.1 62.1
1.7M 66.9 63.5 68.4 76.0 32.2 89.1 82.1 80.8 29.0 93.7 36.0 60.3 -5.8 8.5 44.9 55.0
II Additional Architecture Details
II-A Modality indicator
Modality indicators $\mathcal{I}_{m^{\prime}}$ are learnable tokens that are appended with encoded tokens $\mathbf{p}$ to generate the output tokens $\mathbf{q}$ through the modality-shared decoder. In the decoder, modality indicators $\mathcal{I}_{m^{\prime}}$ have a default shape of $\mathcal{R}^{1\times d_{out}}$, where $d_{out}$ is the dimension of output tokens. For a certain task that requires $N^{\prime}$ output tokens, we repeat the modality indicator $N^{\prime}$ times, resulting in a shape of $\mathcal{R}^{N^{\prime}\times d_{out}}$.
II-A1 Weight sharing.
As an indicator, modality indicators $\mathcal{I}_{m^{\prime}}$ are shared within the same modality but across different tasks, e.g., the dense labeling modality indicator $\mathcal{I}_{D}$ is utilized for both human parsing task and 2D pose estimation task. This design together with the modality-shared encoder and decoder helps Hulk learn human-centric knowledge, remedying task bias from a certain task.
II-A2 Shape of modality indicators.
While the aligning text features $\mathbf{v}\in\mathcal{R}^{C\times d_{out}}$ always have a length of $C$ denoting the number of semantics/classes, $N^{\prime}$ varies among different human-centric tasks. For action and attribute recognition tasks, $N^{\prime}=C$ where $C$ is the number of action or attribute classes. After aligning with BERT semantic features $\mathbf{v}$, the similarity in the $i$-th index represents the existence probability of the $i$-th action/attribute. For the image caption task, $N^{\prime}=40$ means the length of caption tokens. For dense label output tasks, i.e., human parsing and 2D pose estimation tasks, $N^{\prime}$ equals the number of input image tokens $N$, generating predicted semantic maps with the same resolution. For sparse label output tasks, i.e., pedestrian detection, 3D pose estimation, and mesh recovery task, $N^{\prime}$ equals the number of predicted boxes/joints/vertices.
II-B Positional Embeddings
II-B1 Positional embeddings in the encoder
Different from UniHCP [ci2023unihcp], we do not use a learnable positional embedding, instead, we use the bilinear interpolation of fixed MAE [he2022masked] positional embeddings in the encoder.
II-B2 Positional embeddings in the decoder
As the inputs for the decoder contain two parts, i.e., encoded tokens $\mathbf{p}$ and modality indicator $\mathcal{I}_{m^{\prime}}$, therefore, positional embeddings in the decoder have two parts.
For the encoded tokens part, positional embeddings are set as the bilinear interpolation of fixed MAE positional embeddings.
For the modality indicator part, we prepare three types of positional embeddings for better experimental results across diverse human-centric tasks:
  • Anchor points. For the pedestrian detection task, we follow anchor DETR [wang2022anchor] to integrate the location information into modality indicators for achieving better performance. To ensure compatibility with the MAE positional embeddings, which have a shape of $14\times 14$, used in the encoded tokens part, we generate $17\times 17$ positional embeddings as fixed anchor points. This approach enables Hulk to focus on regressing the distances between these anchor points and the actual ground truth boxes. By doing so, the detection task is simplified, effectively reducing the complexity and difficulty of convergence.
  • 3D positional embeddings. For 3D pose estimation and mesh recovery tasks, Hulk needs to predict 3D digit locations using 3D positional embeddings. Following the practice in [feichtenhofer2022masked], we separate the 3D positional embeddings into a fixed $2D$ positional embedding and a fixed $1D$ positional embedding with $PE_{3D}=PE_{2D}+PE_{1D}$, introducing strong 3D location information into 3D regression tasks. The $2D$ positional embedding and $1D$ positional embedding can be generated by interpolating from fixed MAE positional embeddings.
  • Simple interpolation. For tasks encompassing 2D vision, vision-language, and skeleton-based applications, we employ a straightforward approach to generate the positional embeddings in the modality indicator part. These embeddings are inherently two-dimensional or one-dimensional and are derived through a process of simple interpolation from the fixed MAE positional embeddings.
II-C Objective function details
In this section, we present the objective functions in Sec. IV.D in the main text. Given the prediction $\hat{\mathbf{y}}_{m}^{\prime}$ where $m^{\prime}\in\{I,T,S,D\}$ is the output modality, we compute the losses according to different human-centric tasks. As modality-specific de-tokenizers transform output tokens to the semantic part and the digital location part, we adopt two types of losses, i.e., semantic contrastive loss and digit regression loss.
For semantic constrasitve loss, we first compute the conv-transformed output tokens $\hat{\mathbf{f}}\in\mathbb{R}^{N^{\prime}\times d_{out}}$ with $\hat{\mathbf{f}}=\text{Conv}(\mathbf{q})$, where $N^{\prime}$ is the number of output tokens, $d_{out}$ is the output dimension, and $\mathbf{q}\in\mathbb{R}^{N^{\prime}\times d_{out}}$ are output tokens. Given $\mathbf{v}=(\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{C})$ containing the BERT features of $C$ ground truth classes, the semantic contrastive loss is defined as
$\mathcal{L}^{s}=\frac{\mathbf{v}_{k}^{\top}\hat{\mathbf{f}}}{\mathbf{v}_{1}^{\top}\hat{\mathbf{f}}+\mathbf{v}_{2}^{\top}\hat{\mathbf{f}}+...+\mathbf{v}_{C}^{\top}\hat{\mathbf{f}}},$ (1)
where $k$ is the index of the selected semantic token.
For digit regression loss, we compute the distance between predicted coordinate values $\hat{\mathbf{y}}^{d}$ and corresponding ground truth coordinate values $\mathbf{y}^{d}$ with
$\mathcal{L}^{d}=Dis(\hat{\mathbf{y}}^{d},\mathbf{y}^{d}),$ (2)
where $Dis(\mathbf{a},\mathbf{b})$ measures the distance between $\mathbf{a}$ and $\mathbf{b}$, which is different in different tasks, i.e., L1 loss and GIoU loss in the pedestrian detection.
Human Parsing. Human parsing aims to segment human parts. As human parts can be represented by their BERT features, we align the output tokens with BERT features of semantic classes using per-pixel contrastive loss. Following the practice in UniHCP [ci2023unihcp], we also supervise the global probability of a certain semantic class to enhance the performance. Given the conv-transformed output tokens $\hat{\mathbf{f}}$, we simply average pool them to represent the predicted probability of occurrence of certain semantic class $\hat{\mathbf{f}}_{avg}$. Given extracted BERT features $\mathbf{v}$ of $C$ parsing classes, , the total objective function for human parsing is as follows:
$\mathcal{L}_{par}$ $=\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}_{pix})+\mathcal{L}^{s}(\hat{\mathbf{f}}_{avg},\mathbf{v})+\mathcal{L}_{dice}(\hat{\mathbf{y}},y),$ (3)
where $\mathbf{v}_{pix}\in\mathbb{R}^{N\times C\times d_{out}}$ denotes the per-pixel BERT features, $L_{dice}$ is adopted as an auxiliary loss to enhance the human parsing performance. In the dice loss, $y\in\mathbb{R}^{N\times C}$ denotes the ground truth heatmaps with $C$ parsing classes and $\hat{\mathbf{y}}_{D}\in\mathbb{R}^{N\times C}$ is the predicted heatmaps computed by $\hat{\textbf{y}}=\{\hat{y}_{j}=\text{argmax}_{k\in[1,C]}\mathbf{v}^{\top}\text{Upsample}(\hat{\mathbf{f_{j}}})\},j\in[1,N^{\prime}]$.
2D Pose Estimation. We follow the common top-down setting for 2D pose estimation. As we leverage the heatmap-based methods for 2D pose estimation for better experimental results, the objective function is similar to that in human parsing. Mathematically, given extracted BERT features $\mathbf{v}$ of $C$ joint names, we have
$\mathcal{L}_{pose}=\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}_{pix})+\mathcal{L}^{s}(\hat{\mathbf{f}}_{avg},\mathbf{v}),$ (4)
where $\mathbf{v}_{pix}\in\mathbb{R}^{N\times C\times d_{out}}$ denotes the per-pixel BERT features.
Pedestrian Attribute Recognition. Pedestrian Attribute Recognition predicts whether an attribute exists in the input image. Therefore, we only supervise the output tokens $\hat{\mathbf{f}}\in\mathbb{R}^{C\times d_{out}}$ by aligning them to $C$ different attribute semantics. High similarity means the existence of a certain attribute. Given extracted attribute BERT features $\mathbf{v}\in\mathbb{R}^{C\times d_{out}}$, the objective function can be computed by:
$\mathcal{L}_{attr}=\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}).$ (5)
Pedestrian Image Caption. We predict the image caption in an auto-regressive manner. For each predicted token, we adopt the semantic constrastive loss $\mathcal{L}^{s}$ to supervise. The total objective function is similar to that in attribute recognition,
$\mathcal{L}_{cap}=\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}),$ (6)
where $\hat{\mathbf{f}}\in\mathbb{R}^{N^{\prime}\times d_{out}}$ are the features representing predicted caption tokens, $N^{\prime}$ is 40 by default. $\mathbf{v}\in\mathbb{R}^{C\times d_{out}}$ denotes the features of BERT vocabulary, where $C=30522$ denotes the length of BERT vocabulary.
Pedestrian Detection. Different from widely adopted designs in transformer-based pedestrian detectors that rely on object queries, we directly decode $N^{\prime}$ modality indicators as the sparse prediction results. Given predicted digital location $\hat{\mathbf{y}}^{d}$, predicted semantic features $\hat{\mathbf{f}}$, ground truth location $\mathbf{y}^{d}$ and ground truth semantic features $\mathbf{v}$ representing the word “pedestrian”, we utilize optimal bipartite matching to determine the matched pairs. The objective function can be computed by the combination of semantic contrastive loss and digit regression loss:
$\mathcal{L}_{det}=\mathcal{L}_{L1}(\hat{\mathbf{y}}^{d},\mathbf{y}^{d})+\mathcal{L}_{iou}(\hat{\mathbf{y}}^{d},\mathbf{y}^{d})+\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}),$ (7)
where $\mathcal{L}_{L1}$, $\mathcal{L}_{iou}$ are L1 loss and GIoU loss, respectively.
3D Pose Estimation & Mesh Recovery. Following common practice [fastmetro, metro, li2022cliff], we train Hulk to learn 3D pose estimation task and mesh recovery task jointly. Similar to the pedestrian detection task, the 3D pose estimation task also has a combined objective function of semantic contrastive loss and digital regression loss while the mesh recovery task only has the digit regression loss due to undefined semantics of numerous vertices. Given predicted digital location $\hat{\mathbf{y}}^{d}$, predicted semantic features $\hat{\mathbf{f}}$ representing $C$ different joint semantics, we compute the objective function for 3D pose estimation and mesh recovery as
$\mathcal{L}_{3dpos}+\mathcal{L}_{mesh}=$ $\mathcal{L}_{2d}(\hat{\mathbf{y}}^{d},\mathbf{y})+\mathcal{L}_{3d}(\hat{\mathbf{y}}^{d},\mathbf{y})$ (8)
$+\mathcal{L}_{vertice}(\hat{\mathbf{y}}^{d},\mathbf{y})+\mathcal{L}_{normal}(\hat{\mathbf{y}}^{d},\mathbf{y})$
$+\mathcal{L}_{edge}(\hat{\mathbf{y}}^{d},\mathbf{y})+\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}),$
where $\mathbf{y}$ is the ground truth location, $\mathbf{v}$ denotes BERT features of joint names. $\mathcal{L}_{2d}$, $\mathcal{L}_{3d}$, $\mathcal{L}_{vertice}$, $\mathcal{L}_{normal}$, and $\mathcal{L}_{edge}$ are all L1 loss.
Skeleton-based action Recognition. Skeleton-based action recognition predicts the action based on input skeleton sequence. Therefore, given action semantic features $\mathbf{v}\in\mathbb{R}^{C\times d_{out}}$ where $C$ is the number of actions, and the predicted semantic tokens $\hat{\mathbf{f}}\in\mathbb{R}^{C\times d_{out}}$, the objective function can be computed by:
$\mathcal{L}_{ske}=\mathcal{L}^{s}(\hat{\mathbf{f}},\mathbf{v}).$ (9)
TABLE III: Statistics of Training Datasets in the full setting.
Task type Dataset Number of samples
Human3.6M [h36m_pami] 62,668
LIP [lip] 30,462
Human parsing CIHP [cihp] 28,280
ModaNet [zheng2018modanet] 52,245
(7 datasets) VIP [vip] 18,469
Deep fashion [ge2019deepfashion2] 191,961
Paper Doll [paper] 1,035,825
COCO [coco] 149,813
AIC [aic] 378,352
2D pose estimation Human3.6M (pose) [h36m_pami] 312,187
Posetrack [andriluka2018posetrack] 97,174
(8 datasets) 3DPW [3dpw] 68,663
JRDB-Pose [jrdb] 139,385
MPI-INF-3DHP [mehta2017monocular] 1,031,701
AIST++ [aist] 1,015,257
RAPv2 [rapv2] 67,943
Attribute recognition PA_100k [pa] 90,000
Parse27k [PARSE27k] 27,482
(6 datasets) Market [market] 12,926
HARDHC [HARDHC] 28,336
LU-Person [fu2021unsupervised] 10,684,342
CrowdHuman [shao2018crowdhuman] 15,000
Pedestrian detection EuroCity [braun2018eurocity] 21,785
CityPersons [zhang2017citypersons] 2,778
(6 datasets) WiderPerson [zhang2019widerperson] 9,000
WiderPedestrian [loy2019wider] 58,009
COCO-person [coco] 64,115
NTU60 [shahroudy2016ntu] 40,091
Skeleton-based action NTU120 [liu2019ntu] 63,026
GYM [shao2020finegym] 2,778
(6 datasets) Diving48 [diving] 15,027
UCF101 [soomro2012ucf101] 13,320
Kinetics-400 [kay2017kinetics] 239,737
Image caption
CUHK-PEDS [cuhk] 68,126
(2 datasets) SYNTH-PEDES [zuo2023plip] 12,138,157
Human3.6M [h36m_pami] 312,188
3DPW [3dpw] 22,735
3D pose & Mesh COCO [coco] 40,055
MUCO [muco] 101,883
(7 datasets) UP-3D [up3d] 7,126
MPII [mpii] 14,810
GTA [gta] 1,396,913
Total
42 30,187,836$\approx$30M
TABLE IV: Statistics of Training Datasets in the ablation setting.
Task type Dataset Number of samples
Human parsing Human3.6M [h36m_pami] 62,668
LIP [lip] 30,462
(3 datasets) CIHP [cihp] 28,280
2D pose estimation
COCO [coco] 149,813
(2 datasets)
AIC [aic] 378,352
RAPv2 [rapv2] 67,943
Attribute recognition PA_100k [pa] 90,000
Parse27k [PARSE27k] 27,482
(5 datasets) Market [market] 12,926
HARDHC [HARDHC] 28,336
Pedestrian detection
CrowdHuman [shao2018crowdhuman] 15,000
Skeleton action
NTU60 [shahroudy2016ntu] 40,091
(2 datasets)
NTU120 [liu2019ntu] 63,026
Image caption
CUHK-PEDS [cuhk] 68,126
(2 datasets) SYNTH-PEDES [zuo2023plip] 132,000
Human3.6M [h36m_pami] 312,188
3DPW [3dpw] 22,735
3D pose & Mesh COCO [coco] 40,055
MUCO [muco] 101,883
(6 datasets) UP-3D [up3d] 7,126
MPII [mpii] 14,810
Total
21 1,693,302$\approx$1.7M
III Details of training datasets
III-A Full setting
As shown in Table III, we utilize 42 publically available datasets to form the training set for training our Hulk model, containing 30,187,836 training samples and covering eight different human-centric tasks, including human parsing, 2D pose estimation, attribute recognition, pedestrian detection, skeleton-based action recognition, image caption, 3D pose estimation, and mesh recovery.
III-B Ablation setting
In ablation stuides, to demonstrate the effectiveness of Hulk and further explore the potential of a human-centric foundation model, we conduct several ablations on a smaller training set. In Table IV, the ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks.
IV Details of dataset-wise configurations
IV-A Training
We provide detailed dataset-wise training configurations in Table V, including total batch size, batch size per GPU, the number of GPUs, and task weights. In order to better unify different tasks and different datasets, the different datasets of certain tasks (i.e., attribute recognition, pedestrian detection, skeleton-based action, image caption, 3D pose and mesh recovery) are stacked together as a integrated dataset for training, and the datasets of some other tasks (i.e., human parsinga nd 2D pose estimation) are separately trained on different GPUs with diverse task weights.
IV-B Fine-tuning
For fine-tuning porcess, we carefully tune the learning rate, total batch size, the number of iterations, and drop path rate, and report the best performances. In details, the ViT-base backbone (i.e., in Table VI) and ViT-large backbone (i.e., in Table VII) are tuned, respectively.
TABLE V: Hulk joint training setup.
Task Type Dataset Batch Size Batch Size per GPU GPUs Task Weight
Human parsing LIP 108 27 4 1.8
CIHP 104 26 4 3.6
Human3.6M 217 31 7 2.25
ModaNet 27 27 1 0.021
VIP 27 27 1 0.021
Deepfashion 54 27 2 0.042
PaperDoll 27 27 1 0.021
2D pose estimation COCO 528 176 3 28000
AIC 1323 189 7 56000
Human3.6M 264 132 2 3192
Posetrack 340 170 2 12335
JRDB 340 170 2 8223
MPI-INF-3DHP 340 170 2 8223
3DPW 170 170 1 2055
AIST++ 170 170 1 2055
Attribute recognition RAPv2 147 147 1 5
PA
Parse27k
Market
HARDHC
LUPerson 300 300 1 5
Pedestrian detection CrowdHuman 32 4 8 15
EuroCity 80 4 20 42.4
CityPersons
WiderPerson
WiderPedestrian
COCO
Skeleton action Ntu60 240 120 2 4.4
Ntu120
gym
Diving48 90 90 1 1
UCF101
K400
Image caption SYNTH-PEDES 300 100 3 90
CUHK-PEDES
3D pose & Mesh 3DPW 495 165 3 0.5
Human3.6M
COCO
mMUCOuco
UP-3D
MPII
GTA
TABLE VI: Detailed finetuning configs with ViT-base backbone for human-centric tasks.
Task Type Dataset Learning Rate Batch Size Iterations Drop Path Rate
2D pose estimation COCO 3.00E-05 1024 20k 0.3
AIC 3.00E-04 1024 15k 0.2
3D pose & Mesh 3DPW 3.00E-05 512 10k 0.2
Human3.6M 3.00E-05 512 10k 0.2
Pedestrian detection CrowdHuman 1.00E-04 32 80k 0.2
Human parsing LIP 1.00E-04 1024 20k 0.3
CIHP 1.00E-04 1024 20k 0.3
Human3.6M 5.00E-04 1024 20k 0.3
Attribute recognition RAPv2 6.00E-04 128 8k 0.2
PA 3.00E-04 128 5k 0.2
Skeleton action NTU60 1.00E-04 48 10k 0.2
Image caption CUHK-PEDES 1.00E-06 256 3k 0.2
TABLE VII: Detailed finetuning configs with ViT-large backbone for human-centric tasks.
Task Type Dataset Learning Rate Batch Size Iterations Drop Path Rate
2D pose estimation COCO 3.00E-05 1024 10k 0.6
AIC 3.00E-04 1024 15k 0.3
3D pose & Mesh 3DPW 3.00E-05 512 10k 0.2
Human3.6M 3.00E-05 512 10k 0.2
Pedestrian detection CrowdHuman 1.00E-04 32 10k 0.5
Human parsing LIP 5.00E-05 512 10k 0.5
CIHP 5.00E-05 512 10k 0.5
Human3.6M 1.00E-05 512 10k 0.5
Attribute recognition RAPv2 1.00E-03 128 6k 0.5
PA 7.00E-04 128 3k 0.5
Skeleton action NTU60 1.00E-05 48 10k 0.5
Image caption CUHK-PEDES 1.00E-06 256 3k 0.5